OpenStudy (anonymous):

If f(x)=x^2-2x, what is the value of f(y+2)?

OpenStudy (anonymous):

Substitute x=y+2 and reduce the expression. That is (y+2)^2-2*(y+2)

OpenStudy (anonymous):

\[f(y+2)=(y+2)^2-2(y+2)\] then algebra

OpenStudy (anonymous):

\[(y+2)(y+2)-2(y+2)=y^2+4y+4-2y-4=y^2+2y\]

OpenStudy (anonymous):

Answer is f(y+2) is y^2 + 2y. You get this by using his formula. Lol. People type too fast. :)

OpenStudy (anonymous):

thanks everyone!